How do I solve for addition x-2y=-1, -x+5y=4 when I don't have numbers for x.
I tried x-2y=-1 (x=2y+-1) x=-3,
solve for addition x-2y=-1
-x+5y=4

Answer by Alan3354(69443)

How do I solve for addition x-2y=-1, -x+5y=4 when I don't have numbers for x.
I tried x-2y=-1 (x=2y+-1) x=-3,
solve for addition x-2y=-1
-x+5y=4
-------------
I don't know what "solve for addition" means, but
+x-2y=-1
-x+5y=4
---------- Add the equations
3y = 3
y = 1
-------
x-2y = -1
x-2 = -1
x = 1